# Can LLMs Express Personality Across Cultures? Introducing CulturalPersonas for Evaluating Trait Alignment

### v1

@@ Abstract
As LLMs become central to interactive applications, ranging from tutoring to mental health, the ability to express personality in culturally appropriate ways is increasingly important. While recent works have explored personality evaluation of LLMs, they largely overlook the interplay between culture and personality. To address this, we introduce CulturalPersonas, the first large-scale benchmark with human validation for evaluating LLMs’ personality expression in culturally grounded, behaviorally rich contexts. Our dataset spans 3,000 scenario-based questions across six diverse countries, designed to elicit personality through everyday scenarios rooted in local values. We evaluate three LLMs, using both multiple-choice and open-ended response formats. Our results show– CulturalPersonas improves alignment with country-specific human personality distributions (over a 20% reduction in Wasserstein distance across models and countries) and elicits more expressive, culturally coherent outputs compared to existing benchmarks. CulturalPersonas surfaces meaningful modulate trait outputs in response to culturally grounded prompts, offering new directions for aligning LLMs to global norms of behavior. By bridging personality expression and cultural nuance, we envision that CulturalPersonas will pave the way for more socially intelligent and globally adaptive LLMs. Datasets and code are available at: https://github.com/limenlp/CulturalPersonas.
@@
Priyanka Dey  Yugal Khanter  Aayush Bothra  Jieyu Zhao  Emilio Ferrara Department of Computer Science, University of Southern California {deyp, khanter, abothra, jieyuz, emiliofe}@usc.edu
@@ 1 Introduction
While these tools offer a useful starting point, they fall short of capturing how personality is expressed through behavior in real-world contexts. Moreover, they assume that personality is culturally universal despite decades of work in cultural psychology, which shows that trait expression is deeply shaped by local norms, values, and communicative styles Hofstede (2001); Han et al. (2023); Liu et al. (2024). For example, assertiveness may reflect Openness in the United States but be seen as disruptive in Japan, where harmony and indirectness are preferred. Evaluating LLMs without accounting for these distinctions risks reinforcing Western behavioral norms and misrepresenting how personality is interpreted across cultures.
@@
We use CulturalPersonas to evaluate three models— GPT-4o-mini, Llama-3-8B, and Qwen2-7B— across two evaluation formats: multiple-choice selection and open-ended generation. Our contributions are fourfold:
@@
  • Dual-Format Evaluation Framework: We propose a two-part evaluation: multiple-choice selection (MCS) and open-ended generation (OEG)—that captures both trait alignment and expressive diversity. This framework enables both distributional comparison with human norms and stylistic analysis of personality expression in various contexts (§4).
@@ 2.1 Personality Evaluation Methods
Table 1: Pairwise Wasserstein distances ($W$) between selected country pairs for each OCEAN trait. Statistical significance is measured via 2-sample KS test (*p $<$ 0.05, **p $<$ 0.01). Larger $W$ indicates higher cross-cultural differences.
@@ 3.3 Human Validation
Figure 2: Inter-annotator agreement (IAA) scores across validation tasks
@@ 3.3.3 Q&A Generation
Since personality expression depends on both situational context and response variability, we next evaluate the quality of question–answer (Q&A) pairs. Annotators assess whether each question logically follows from the scenario and whether the Likert-scale options span a realistic range of trait expression. High agreement scores confirm the benchmark’s coherence and trait sensitivity.
@@ 4 Experimental Setup
Figure 3: $W$ and $KS$ metrics (average over OCEAN dimensions) for the MCS setting across all personality evaluators for GPT-4o-mini. Lower values indicate closer alignment with human distributions. Across all six benchmarked countries, CulturalPersonas yields the closest personality alignment.
@@ 4.3 Generating Model Trait Distributions
To simulate culturally contextualized behavior, we use cultural priming by providing each model with a light persona prompt, inspired from previous works Li et al. (2024a); Kirk et al. (2024): “Imagine you are someone born and raised in {country}. You are very familiar with {country}’s culture and traditions and practice many of them, two of which include: {norm1} and {norm2}.” Norms are drawn from each country-specific knowledge base retrieved via our RAG pipeline (§ 3).
@@ 4.3.2 Open-Ended Generation (OEG)
As there is no gold-standard Likert score for model-generated outputs, we adopt a pseudo-labeling strategy to estimate trait scores. We first embed each response using sentence-transformers Reimers and Gurevych (2019) and then compute cosine similarities between the response and each of the predefined answer options for that item (2 for Big5Chat, 4 for TRAIT, 5 for CulturalPersonas). These similarity scores form a soft probability distribution over the 1–5 Likert scale. To obtain a final trait score distribution per model response, we sample $n$ times (total human samples from ground truth) from this soft distribution. We validate the effectiveness of this approach via a human study in which annotators match model responses to the predefined answer sets (§5.1).
@@ 5.1 Alignment with Human Personalities
Table 2 summarizes relative gains in $W$ using CulturalPersonas over all personality evaluators averaged over models. We find robust improvements across countries and baselines, up to +25.9% in South Africa (vs. BFI) and +32.8% in Japan (vs. Big5Chat). These gains are benchmark-agnostic, suggesting that our culturally grounded scenarios drive stronger alignment.
[…]
@@ 5.3 Model Variation in Personality Expression
Figure 5: $W$ across OCEAN traits and models (avgerage over countries) for MCS. GPT-4o-mini performs best overall. CulturalPersonas reveals models struggle with traits: Agreeableness and Extraversion across cultures.
@@ 5.4 Role of Cultural Priming in Trait Expression
Finally, we isolate the impact of cultural priming via an ablation study. We remove cultural cues from prompts—excluding the country name and norm preamble—while retaining the scenario and question specification. As shown in Figure 6, this significantly degrades alignment: average $W$s in Brazil rise by 31.2%. Traits like Agreeableness and Extraversion show the largest drops, with models reverting to neutral, moderate expressive language.
For example, in a team feedback scenario—“A colleague has just presented a project. You are asked to provide feedback.”—the unprimed model gives a direct and neutral response: “Your points were good, but I don’t think this section should be here. Instead, you should move it to when you are discussing x.” However, when primed with information about Japan and its norms of collective harmony and group cohesion, the same model instead says: “Thank you for your presentation. If you’d like, I have some feedback to improve your flow. We can discuss privately.” [^3] More examples included in Appendix.
Figure 6: Effect of cultural priming for Brazil across models and personality evaluators. Priming consistently reduces deviation across benchmarks. Bars show relative improvement in $W$ (%); * and ** denote $p<0.05$ and $p<0.01$ (Wilcoxon signed-rank test over 5 traits).
@@ 6 Conclusion & Future Work
7 Limitations
@@
Finally, our evaluation captures trait expression at a single point in time, within static scenarios. This snapshot-style analysis does not reflect whether LLMs can maintain consistent persona behavior over time or across diverse interactions—a key aspect of psychological personality. Studying temporal consistency and contextual adaptability remains an important direction for future exploration.
8 Ethics Statement
@@ Appendix D Additional Results
In this section, we present additional analyses and results for personality expression in LLMs. In particular, we focus on (1) Alignment metrics for all 3 models across all 6 benchmarked countries (§ D.1), (2) Relative improvements in personality alignment with trait breakdowns (§ D.1), (3) Open-ended personality expression across various benchmarks and models (§ D.2), (4) Ablation results from each country (§ E).
@@ D.1 Personality Alignment Across Various LLMs
We first present all results for multiple choice selection (MCS) and open-ended generation (OEG) split by model in Figures 10 and 11. Our results show that MCS setting is much easier with lower alignment metrics. However, in both settings, CulturalPersonas can yield closer alignment with real human populations. GPT-4o-mini is able to achieve the strongest alignment across most countries. Qwen2 however tends to struggle more with expressive and underrepresented countries such as Brazil and South Africa. Llama3 tends to perform on average (in the middle) for most countries, only beating GPT-4o-mini on India.
@@
Next, we present the relative improvement rates for each country using CulturalPersonas over both settings: MCS and OEG in Tables 6 and 7. We find that the models generally show strong yields for Conscientiousness and Openness, with more moderate to low gains for traits such as Neuroticism and Agreeableness. In particular, gains are most pronounced in East Asian countries i.e. Japan, India, and Saudi Arabia and Africa i.e. South Africa.
@@ D.2 Analysis of Open-Ended Generation Results
Additionally, we show examples of personality expression across various models. In Tables 9 and 10, we show variations of how Extraversion and Agreeableness appear in Qwen2 v. Llama3 v. GPT-4o-mini. Qwen2 often struggles with expressing Agreeableness in countries e.g. Japan. The example highlights more Western-centric norms in its answers compared to GPT-4o-mini and Llama3 whose responses highlight more of a passive and indirect method for maintaining harmony. On the other hand, Llama3 struggles with expressing Extraversion in expressive nations such as Brazil. We can see when comparing with model responses from Qwen2 and GPT-4o-mini, these responses are much more expressive and engaging in ways related to Brazilian norms.
@@ Appendix E Additional Ablations
Figure 12: Results from cultural priming ablations for each country and model for both MCS and OEG settings. We observe large increases in alignment (Wasserstein distances) for various countries including underrepresented and highly expressive nations.

### v2

@@ Abstract
As LLMs become central to interactive applications, ranging from tutoring to mental health, the ability to express personality in culturally appropriate ways is increasingly important. While recent works have explored personality evaluation of LLMs, they largely overlook the interplay between culture and personality. To address this, we introduce CulturalPersonas, the first large-scale benchmark with human validation for evaluating LLMs’ personality expression in culturally grounded, behaviorally rich contexts. Our dataset spans 3,000 scenario-based questions across six diverse countries, designed to elicit personality through everyday scenarios rooted in local values. We evaluate how closely three models’ personality distributions align to real human populations through two evaluation settings: multiple-choice and open-ended response formats. Our results show– CulturalPersonas improves alignment with country-specific human personality distributions (over a 20% reduction in Wasserstein distance across models and countries) and elicits more expressive, culturally coherent outputs compared to existing benchmarks. CulturalPersonas surfaces meaningful modulate trait outputs in response to culturally grounded prompts, offering new directions for aligning LLMs to global norms of behavior. By bridging personality expression and cultural nuance, we envision that CulturalPersonas will pave the way for more socially intelligent and globally adaptive LLMs. Datasets and code are available at: https://github.com/limenlp/CulturalPersonas.
@@
Priyanka Dey   Yugal Khanter   Aayush Bothra Jieyu ZhaoEmilio Ferrara Department of Computer Science, University of Southern California deyp@usc.edu
@@ 1 Introduction
While these tools offer a useful starting point, they fall short of capturing how personality is expressed through behavior in real-world contexts. Moreover, they assume that personality is culturally universal despite decades of work in cultural psychology, which shows that trait expression is deeply shaped by local norms, values, and communicative styles Hofstede (2001); Han et al. (2023); Liu et al. (2024). For example, in Western cultures such as the United States, assertiveness may reflect high openness. However, in nations such as Japan, this same quality may be seen as disruptive; instead, traits such as harmony and indirectness are stronger and more preferred indicators of Openness. Evaluating LLMs without accounting for these distinctions risks reinforcing Western behavioral norms and misrepresenting how personality is interpreted across cultures.
@@
We use CulturalPersonas to measure the personality profiles of three models— GPT-4o-mini, Llama-3-8B, and Qwen2-7B— across two evaluation formats: multiple-choice selection and open-ended generation and compare alignment to real human populations from diverse cultural backgrounds. Our contributions are fourfold:
@@
  • Dual-Format Evaluation Framework: We propose a two-part evaluation: multiple-choice selection (MCS) and open-ended generation (OEG)—that captures both trait alignment with real user populations and expressive diversity. This framework enables both distributional comparison with human norms and stylistic analysis of personality expression in various contexts (§4).
@@ 2.1 Personality Evaluation Methods
Table 1: Pairwise Wasserstein distances ($W$) between selected country pairs for each OCEAN trait. Statistical significance is measured via 2-sample KS test (*p << 0.05, **p << 0.01). Larger $W$ indicates higher cross-cultural differences.
@@ 3.3 Human Validation
Country
Norm
Relevance
Scenario
Quality
Q&A
Generation
Brazil 1.00 0.98 0.92
India 1.00 0.98 0.96
USA 1.00 0.98 0.99
Japan 1.00 0.96 0.98
Saudi Arabia 0.90 0.93 0.94
South Africa 1.00 0.94 0.96
Table 2: Inter-annotator agreement (IAA) scores across human validation tasks for CulturalPersonas generation.
@@ 3.3.3 Q&A Generation
Since personality expression depends on both situational context and response variability, we next evaluate the quality of question–answer (Q&A) pairs. Annotators assess whether each question logically follows from the scenario and whether the Likert-scale options span a realistic range of trait expression. High agreement scores (IAA $\geq$ 0.9) show the diversity and variance in questions in CulturalPersonas.
@@ 4 Experimental Setup
Figure 2: Wasserstein distance and KS statistic metrics (average over OCEAN dimensions) for the MCS setting across all personality evaluators for GPT-4o-mini. Lower values indicate closer alignment with human distributions. Across all six benchmarked countries, CulturalPersonas yields the closest personality alignment.
@@ 4.3 Generating Model Trait Distributions
To simulate culturally contextualized behavior, we use cultural priming by providing each model with a light persona prompt, inspired from previous works Li et al. (2024a); Kirk et al. (2024): “Imagine you are someone born and raised in {country}. You are very familiar with {country}’s culture and traditions and practice many of them, two of which include: {norm1} and {norm2}.”
Country-specific norms are extracted from each knowledge base via our RAG pipeline (§ 3). We retain the top two norms for which all annotators concurred, i.e., those consistently deemed both relevant and salient for the respective country. Although inclusion of all ten norms was feasible, we adhere to prior work Liu et al. (2023); Leng et al. (2024) indicating that excessively long prompts or contextual information can impair model performance.
@@ 4.3.2 Open-Ended Generation (OEG)
Since no gold-standard Likert scores exist for model free responses, we adopt a pseudo-labeling strategy. Given a model response $r$ and predefined answer options $A=\{a_{1},\dots,a_{k}\}$ [^3] The value of $k$ varies based on personality evaluator. For BFI, IPIP-120, and IPIP-300, $k=5$; for TRAIT, $k=4$, for Big5Chat, $k=2$., we embed both $r$ and each $a_{i}$ using sentence-transformers Reimers and Gurevych (2019). We then compute cosine similarities:
$s_{i}=\cos(E(r),E(a_{i})),\quad i=1,\dots,k,$
and then normalize the set of similarities into a soft distribution:
$p_{i}=\frac{\exp(s_{i})}{\sum_{j=1}^{k}\exp(s_{j})}.$
Each option $a_{i}$ maps to a Likert score $\ell_{i}\in\{1,\dots,5\}$, producing a trait distribution:
$P(\ell)=\sum_{i=1}^{k}p_{i}\,\delta(\ell=\ell_{i}).$
Finally, to generate model score distributions for each Big5 trait, we draw $n$ samples from this distribution, where $n$ matches the number of ground-truth annotations. We validate the effectiveness of this approach via a human study where each annotator matches a set of model personality evaluator responses to the predefined answer sets (§5.1).
@@ 5 Evaluating Multi-Cultural Personalities in LLMs
5.1 Alignment with Human Personalities
Table 3 summarizes relative gains in Wasserstein distances using CulturalPersonas over all personality evaluators averaged over models. We find robust improvements across countries and baselines, up to +25.9% in South Africa (vs. BFI) and +32.8% in Japan (vs. Big5Chat). These gains are benchmark-agnostic, suggesting that our culturally grounded scenarios drive stronger alignment.
We observe a similar trend in the OEG setting, where models generate free-form text in response to scenario prompts. To assign trait scores, we embed generations using sentence-transformers and compute similarity to Likert-style trait options. To validate this scoring pipeline, we conduct a human annotation study (Figure 3). Annotators rated 540 model generations from the OEG setting (10 per country for each model and benchmark) on a 1–5 Likert scale based on perceived expression of each trait. We compute Pearson correlations between the auto-mapped and human scores, which yield strong alignment across countries (r=0.61r=0.610.730.73).
@@ 5.2 Linguistic Expressivity in OEG
Figure 4: Wasserstein distances across OCEAN traits and models (avgerage over countries) for MCS. GPT-4o-mini performs best overall. CulturalPersonas reveals models struggle with traits: Agreeableness and Extraversion across cultures.
@@ 5.4 Role of Cultural Priming in Trait Expression
In our experiments, models are given a persona prompt with country information and cultural norms (see §4.3) to guide their interpretation of personality questions. Simpler prompts such as “You are someone from {country} were also tested, but yielded little improvement (Appendix E.2), motivating our use of richer persona descriptions. To isolate the effect of this cultural priming, we compare against a setting where models receive only the personality questions from CulturalPersonas, without the contextual prompt.
As shown in Figure 5, removal of cultural priming significantly degrades alignment: average Wasserstein distances in Brazil rise by 31.2%. Traits like Agreeableness and Extraversion show the largest drops, with models reverting to neutral, moderate expressive language.
For example, in a team feedback scenario—“A colleague has just presented a project. You are asked to provide feedback.”—the unprimed model gives a direct and neutral response in OEG setting: “Your points were good, but I don’t think this section should be here. Instead, you should move it to when you are discussing x.” However, when primed with information about Japan and its norms of collective harmony and group cohesion, the same model instead says: “Thank you for your presentation. If you’d like, I have some feedback to improve your flow. We can discuss privately.” [^4] More examples included in Appendix.
Figure 5: Effect of cultural priming for Brazil across models and personality evaluators. Priming consistently reduces deviation across benchmarks. Bars show relative improvement in Wasserstein distances (%); * and ** denote $p<0.05$ and $p<0.01$ (Wilcoxon signed-rank test over 5 traits).
@@ 6 Conclusion & Future Work
Acknowledgments
This research was supported by the NSF, under Award Number 2331722. The authors of this work would like to thank the support of researchers from USC NLP and the Information Sciences Institute for their continous feedback in the formulation and setting of our work. We would also like to thank the OpenReview ARR anonymous reviewers and area chairs for their helpful suggestions in improving our work.
Limitations
@@ Acknowledgments
Our evaluation captures trait expression at a single point in time, within static scenarios. This snapshot-style analysis does not reflect whether LLMs can maintain consistent persona behavior over time or across diverse interactions—a key aspect of psychological personality. Studying temporal consistency and contextual adaptability remains an important direction for future exploration.
Imbalanced Country Datasets:
Finally, our experiments include ground truth datasets called from the IPIP-NEO data repository which contains varying data points for each country. Due to survey availability in each country, the amount of human ground truth data i.e. the size of personality profiles for each country greatly varies. For instance, for Saudi Arabia, we are only able to compare to a size of 98 users. Although this is a significantly smaller set compared to data from i.e. USA (over 200k users) or India (around 3k users), prior work in cross-cultural psychology Mõttus et al. (2012); Nezlek et al. (2011) has shown that even with as small as 100 users, meaningful conclusions can be drawn from users in varying demographics.
Ethics Statement
@@ Appendix D Additional Results
In this section, we present additional analyses and results for personality expression in LLMs. In particular, we focus on (1) Alignment metrics for all 3 models across all 6 benchmarked countries (§ D.1), (2) Relative improvements in personality alignment with trait breakdowns (§ D.1), (3) Open-ended personality expression across various benchmarks and models (§ D.2).
@@ D.1 Personality Alignment Across Various LLMs
We first present all results for multiple choice selection (MCS) and open-ended generation (OEG) split by model in Figures 9 and 10. Our results show that MCS setting is much easier with lower alignment metrics. However, in both settings, CulturalPersonas can yield closer alignment with real human populations. GPT-4o-mini is able to achieve the strongest alignment across most countries. Qwen2 however tends to struggle more with expressive and underrepresented countries such as Brazil and South Africa. Llama3 tends to perform on average (in the middle) for most countries, only beating GPT-4o-mini on India.
@@
Next, we present the relative improvement rates for each country using CulturalPersonas over both settings: MCS and OEG in Tables 7 and 8. We find that the models generally show strong yields for Conscientiousness and Openness, with more moderate to low gains for traits such as Neuroticism and Agreeableness. In particular, gains are most pronounced in East Asian countries i.e. Japan, India, and Saudi Arabia and Africa i.e. South Africa.
@@ D.2 Analysis of Open-Ended Generation Results
Additionally, we show examples of personality expression across various models. In Tables 10 and 11, we show variations of how Extraversion and Agreeableness appear in Qwen2 v. Llama3 v. GPT-4o-mini. Qwen2 often struggles with expressing Agreeableness in countries e.g. Japan. The example highlights more Western-centric norms in its answers compared to GPT-4o-mini and Llama3 whose responses highlight more of a passive and indirect method for maintaining harmony. On the other hand, Llama3 struggles with expressing Extraversion in expressive nations such as Brazil. We can see when comparing with model responses from Qwen2 and GPT-4o-mini, these responses are much more expressive and engaging in ways related to Brazilian norms.
@@ Appendix E Additional Ablations
In this section, we present ablation experiments by removing our selected persona prompts i.e. removing cultural priming and observing distributional alignment changes between models and human data § E.1. We also present results from using a simple prompt with only a country persona in § E.2.
E.1 Cultural Priming Ablations
@@
Figure 11: Results from cultural priming ablations for each country and model for both MCS and OEG settings. We observe large increases (%) in alignment (Wasserstein distances) for various countries including underrepresented and highly expressive nations.
E.2 Simple Prompting
In this section, we provide results from our simple prompting experiments. Building on prior findings that simple persona prompts often lack sufficient context to elicit demographic-specific responses Tao et al. (2024); Ghosh et al. (2024), we examine the extent to which cultural priming offers more effective guidance. In Table 12, we present the average Wasserstein distances for each country given a simple persona prompt i.e. “You are someone from {country}."
Country BFI IPIP-120 IPIP-300 TRAIT Big5Chat CulturalPersonas
Brazil 19.12 24.52 20.20 26.12 22.02 21.44
India 14.80 16.00 14.62 26.80 23.70 19.23
Japan 16.04 22.98 27.70 34.02 32.80 28.34
USA 12.20 17.22 18.30 16.78 26.24 31.34
Saudi Arabia 21.12 25.02 23.62 24.30 24.40 12.45
South Africa 25.88 22.60 19.82 30.02 32.92 20.23
Table 12: Average relative improvement (%) in Wasserstein distance averaged across OCEAN traits using cultural priming over standard country context prompt for each country and evaluation set.